# Scale-Adaptive Tokens

### v2

@@ Abstract
We propose a one-stage framework for real-time multi-person 3D human mesh estimation from a single RGB image. While current one-stage methods, which follow a DETR-style pipeline, achieve state-of-the-art (SOTA) performance with high-resolution inputs, we observe that this particularly benefits the estimation of individuals in smaller scales of the image (e.g., those far from the camera), but at the cost of significantly increased computation overhead. To address this, we introduce scale-adaptive tokens that are dynamically adjusted based on the relative scale of each individual in the image within the DETR framework. Specifically, individuals in smaller scales are processed at higher resolutions, larger ones at lower resolutions, and background regions are further distilled. These scale-adaptive tokens more efficiently encode the image features, facilitating subsequent decoding to regress the human mesh, while allowing the model to allocate computational resources more effectively and focus on more challenging cases. Experiments show that our method preserves the accuracy benefits of high-resolution processing while substantially reducing computational cost, achieving real-time inference with performance comparable to SOTA methods. Code and models are available at https://ChiSu001.github.io/SAT-HMR/.
Figure 1: (a) We propose scale-adaptive tokens in our one-stage framework for real-time multi-person 3D mesh estimation. Our method introduces scale-adaptive tokens, dynamically adjusted based on the relative size of individuals in the image, to more efficiently encode features, enabling real-time and accurate multi-person mesh estimation. We present a conceptual visualization of the scale-adaptive tokens. The right column visualizes the predicted meshes projected onto an image from 3DPW [48] dataset and from an elevated view. (b) Comparison of estimation error and inference time across different methods, with input resolutions in parentheses. Our method, using a mixed resolution with a base resolution of 644, achieves comparable performance to state-of-the-art methods on AGORA [32] test set while maintaining real-time inference efficiency.
@@ 1 Introduction
Building on the findings of [43, 2], we observe that high-resolution input plays a crucial role in improving accuracy. To explore this, we design a naive baseline following similar architecture [2] and analyze how its estimation error changes for individuals of different scales as we increase the input resolution. We define scale as the size of the person relative to the entire image [^1] See Eq. 1 for detailed definition of scale.. As shown in Fig. 2, increasing the input resolution consistently leads to a decrease in the estimation error, particularly for smaller-scale individuals (e.g., those farther from the camera). Notably, for individuals in the 0-10% scale, we observe an error decrease of nearly 35mm when comparing the highest-resolution model (blueviolet) to the lowest-resolution one (yellow). However, processing the entire image at high resolution incurs significant computational costs, especially when individuals are close to the camera and occupy a large portion of the image. In such cases, these individuals are already represented by many tokens, and using more tokens with high resolution would only greatly increase the cost while offering marginal improvement (i.e., a slight error reduction in cases where individuals cover 30+of the image in Fig. 2). Therefore, both our best baseline model (blueviolet) and recent SOTA methods [43, 2] in Fig. 1 (b) fail to achieve real-time inference and run at about 5 FPS, highlighting a common limitation.
@@ 2.1 Multi-Person Human Mesh Estimation
Multi-person human mesh estimation from a single RGB image aims to regress all human meshes, typically represented by parametric models such as SMPL [24] and SMPL-X [33]. Existing methods can be categorized into multi-stage and one-stage approaches. Multi-stage methods [11, 7, 34, 54, 53, 14] use off-the-shelf human detectors [37] to crop each person’s region, followed by single-person mesh estimation [55, 26, 57, 49]. This approach preserves relatively high-resolution images for each cropped region, generally leading to higher accuracy [43]. However, its reliance on detected bounding boxes and the lack of global context make it challenging to handle occlusion [44].
@@ 3.2 Overview
Naive baseline. While high-resolution inputs are shown to significantly enhance estimation accuracy [43, 2], they also introduce considerable computational overhead. To further explore the trade-off between resolution and efficiency, we first introduce a simple baseline using a similar architecture to [2], as depicted in Fig. 3 (top).
@@ 4.2 Implementation Details
The input image is resized to maintain its original aspect ratio, with a base resolution of $\max(H,W)=644$ and $\max(H_{\text{hr}},W_{\text{hr}})=1288$ for high-resolution when needed. We denote this mixed-resolution setting as “644”. The Transformer layers in our encoder are initialized using pretrained DINOv2 [31] with ViT-B architecture and a patch size of $P=14$. The Transformer used in the encoder has $N_{\text{lr}}=N_{\text{hr}}=3$, and $N_{\text{sa}}=9$ layers, respectively. Our decoder consists of 6 Transformer layers and 50 human queries. Training is conducted on 8 RTX 3090 GPUs with a total batch size of 40 for 60 epochs with an initial learning rate of $54e-$ 5. Please refer to Sec. A.2 for more details.
@@ 4.4 Ablation Study
Background tokens $\mathcal{T}_{\text{B}}$. We investigate the impact of background tokens on model performance and report the MVE across different scale ranges on the BEDLAM validation set in Tab. 4. In (a), we discard all background tokens and observe stable performance at the smallest scale, but a decline as the scale increases, indicating that background context becomes more important for larger scales. In (b), skipping pooling for background tokens yields nearly the same performance but introduces higher computational cost. In (c), applying pooling twice with the same grouping strategy results in a performance drop for large-scale instances, highlighting the significance of background context for estimation. We provide additional efficiency comparisons in Sec. B, showing that pooling once achieves the best balance between performance and efficiency.
@@ B.1 Ablation Study
Speed-accuracy trade-off. We evaluate the speed-accuracy trade-off across various ablation models. Specifically, we compare single-resolution baselines with our model, which adopts different processing strategies for background tokens ($\mathcal{T}_{\text{B}}$). In Tab. B1, we report the average number of tokens, inference runtime, and the MVE metric on the BEDLAM validation set [3]. The average number of tokens is included to highlight the impact of image tokens on inference speed.
@@
Table B1: Speed-accuracy trade-off among different ablation settings. We conduct our studies on BEDLAM [3] validation set, reporting the average number of different tokens, inference runtime and MVE.
 Ablation Number of tokens Runtime (ms) MVE $\downarrow$
High-resolution Low-resolution Background
Single Resolution
(a) Res. 1288 4784 174.9 53.2
(b) Res. 644 1196 42.3 63.3
Background Tokens $\mathcal{T}_{\text{B}}$
(c) Drop all 493 40 0 39.7 57.2
(d) No pooling 493 1073 54.1 56.1
(e) Pooling×$\times$2 493 94 100 41.6 56.3
(f) Ours 493 94 245 42.0 56.0
Table B2: Comparison of different scale thresholds on BEDLAM [3] validation set. We report MVE for different scale ranges and the average (Avg.).
 $\alpha_{\text{s}}$ 0-30% 30-50% 50-70% 70%+ Avg.
(a) 0.0 67.0 60.4 58.6 64.0 64.0
(b) 0.3 60.0 61.9 59.9 65.8 60.8
(c) 0.5 (Ours) 55.6 56.0 57.6 63.1 56.0
(d) 0.7 55.9 56.1 58.2 65.2 56.2
@@
Scale threshold $\alpha_{\text{s}}$. To further study the impact of high-resolution tokens, we conduct experiments on various scale thresholds $\alpha_{\text{s}}$ while retaining the pooling of background tokens. $\alpha_{\text{s}}=0$ denotes that no high-resolution tokens are used. Results are shown in Tab. B2. Compared to (a), (c) shows that our method achieves a consistent error reduction across different scale ranges, indicating that introducing sufficient high-resolution tokens eases the estimation challenge on small-scale instances and also allows the model to deal better with large-scale instances. In (b), although the improvement in the scale range of 0-30% is significant, the decrease in high-resolution samples increases learning difficulty during training and potentially leads to worse performance on larger-scale instances than (a). In (d), a large scale threshold ignores plenty of background context since the high-resolution tokens are encoded independently for $N_{\text{hr}}$ layers, leading to a performance decline, which is consistent with our findings in Sec. 4.4.

### v3

@@ Abstract
We propose SAT-HMR, a one-stage framework for real-time multi-person 3D human mesh estimation from a single RGB image. While current one-stage methods, which follow a DETR-style pipeline, achieve state-of-the-art (SOTA) performance with high-resolution inputs, we observe that this particularly benefits the estimation of individuals in smaller scales of the image (e.g., those of young age or far from the camera), but at the cost of significantly increased computation overhead. To address this, we introduce scale-adaptive tokens that are dynamically adjusted based on the relative scale of each individual in the image within the DETR framework. Specifically, individuals in smaller scales are processed at higher resolutions, larger ones at lower resolutions, and background regions are further distilled. These scale-adaptive tokens more efficiently encode the image features, facilitating subsequent decoding to regress the human mesh, while allowing the model to allocate computational resources more effectively and focus on more challenging cases. Experiments show that our method preserves the accuracy benefits of high-resolution processing while substantially reducing computational cost, achieving real-time inference with performance comparable to SOTA methods.
Figure 1: (a) We propose scale-adaptive tokens in our one-stage framework for real-time multi-person 3D mesh estimation. Our method introduces scale-adaptive tokens, dynamically adjusted based on the relative size of individuals in the image, to more efficiently encode features, enabling real-time and accurate multi-person mesh estimation. We present a conceptual visualization of the scale-adaptive tokens. The right column visualizes the predicted meshes projected onto an image from 3DPW [49] dataset and from an elevated view. (b) Comparison of estimation error and inference time across different methods, with input resolutions in parentheses. Our method, using a mixed resolution with a base resolution of 644, achieves comparable performance to state-of-the-art methods on AGORA [33] test set while maintaining real-time inference efficiency. Code and models are available at https://ChiSu001.github.io/SAT-HMR/.
@@ 1 Introduction
Building on the findings of [44, 2], we observe that high-resolution input plays a crucial role in improving accuracy. To explore this, we design a naive baseline following similar architecture and analyze how its estimation error changes for individuals of different scales as we increase the input resolution. We define scale as the size of the person relative to the entire image [^1] See Eq. 1 for detailed definition of scale.. As shown in Fig. 2, increasing the input resolution consistently leads to a decrease in the estimation error, particularly for smaller-scale individuals (e.g., kids, those far from the camera or those in closed-poses). Notably, for individuals in the 0-10% scale, we observe an error decrease of nearly 35mm when comparing the highest-resolution model (blueviolet) to the lowest-resolution one (yellow). However, processing the entire image at high resolution incurs significant computational costs, especially when individuals are close to the camera and occupy a large portion of the image. In such cases, these individuals are already represented by many tokens, and using more tokens with high resolution would only greatly increase the cost while offering marginal improvement (i.e., a slight error reduction in cases where individuals cover 30+of the image in Fig. 2). Therefore, both our best baseline model (blueviolet) and recent SOTA methods [44, 2] in Fig. 1 (b) fail to achieve real-time inference and run at about 5 FPS, highlighting a common limitation.
@@ 2.1 Multi-Person Human Mesh Estimation
Multi-person human mesh estimation from a single RGB image aims to regress all human meshes, typically represented by parametric models such as SMPL [24] and SMPL-X [34]. Existing methods can be categorized into multi-stage and one-stage approaches. Multi-stage methods [11, 7, 35, 56, 55, 14, 53] use off-the-shelf human detectors [38] to crop each person’s region, followed by single-person mesh estimation [57, 26, 59, 50, 27]. This approach preserves relatively high-resolution images for each cropped region, generally leading to higher accuracy [44]. However, its reliance on detected bounding boxes and the lack of global context make it challenging to handle occlusion [45].
@@ 3.2 Overview
Naive baseline. While high-resolution inputs are shown to significantly enhance estimation accuracy [44, 2], they also introduce considerable computational overhead. To further explore the trade-off between resolution and efficiency, we first introduce a simple baseline using a DETR-style architecture, as depicted in Fig. 3 (top).
@@ 4.2 Implementation Details
The input image is resized to maintain its original aspect ratio, with a base resolution of $\max(H,W)=644$ and $\max(H_{\text{hr}},W_{\text{hr}})=1288$ for high-resolution when needed. We denote this mixed-resolution setting as “644”. The Transformer layers in our encoder are initialized using pretrained DINOv2 [32] with a patch size of $P=14$. We use ViT-B architecture as ViT-L significantly increases inference time (42.0 ms vs. 74.2 ms). The Transformer used in the encoder has $N_{\text{lr}}=N_{\text{hr}}=3$, and $N_{\text{sa}}=9$ layers, respectively. Our decoder consists of 6 Transformer layers and 50 human queries. Training is conducted on 8 RTX 3090 GPUs with a total batch size of 40 for 60 epochs with an initial learning rate of $54e-$ 5. Please refer to Sec. A.2 for more details.
@@ 4.4 Ablation Study
Background tokens $\mathcal{T}_{\text{B}}$. We investigate the impact of background tokens on model performance and report the MVE across different scale ranges on the BEDLAM validation set in Tab. 4. In (a), we discard all background tokens and observe stable performance at the smallest scale, but a decline as the scale increases, indicating that background context becomes more important for larger scales. In (b), skipping pooling for background tokens yields nearly the same performance but introduces higher computational cost. In (c), applying pooling twice with the same grouping strategy results in a performance drop for large-scale instances, highlighting the significance of background context for estimation. We hypothesize that large-scale people face more truncation, relying more on global context which requires further exploration. We provide additional efficiency comparisons in Sec. B, showing that pooling once achieves the best balance between performance and efficiency.
@@ 5 Conclusion
Acknowledgments
This work was supported by 2022ZD0114904 and NSFC-6247070125. We thank Yiyang Wang for the valuable discussions during the early stage of this work.
@@ A.1 Model Architecture
Table A1: Speed-accuracy trade-off among different ablation settings. We conduct our studies on BEDLAM [3] validation set, reporting the average number of different tokens, inference runtime and MVE.
 Ablation Number of tokens Runtime (ms) MVE $\downarrow$
High-resolution Low-resolution Background
Single Resolution
(a) Res. 1288 4784 174.9 53.2
(b) Res. 644 1196 42.3 63.3
Background Tokens $\mathcal{T}_{\text{B}}$
(c) Drop all 493 40 0 39.7 57.2
(d) No pooling 493 1073 54.1 56.1
(e) Pooling×$\times$2 493 94 100 41.6 56.3
(f) Ours 493 94 245 42.0 56.0
@@ B.1 Ablation Study
Speed-accuracy trade-off. We evaluate the speed-accuracy trade-off across various ablation models. Specifically, we compare single-resolution baselines with our model, which adopts different processing strategies for background tokens ($\mathcal{T}_{\text{B}}$). In Tab. A1, we report the average number of tokens, inference runtime, and the MVE metric on BEDLAM [3] validation set. The average number of tokens is included to highlight the impact of image tokens on inference speed.
@@
Table B2: Effect of resolution. We study the impact of resolution using both single-resolution (baseline) and mixed-resolution (our method, with “*”) settings on AGORA [33] and BEDLAM [3], reporting MVE for different scale ranges and the average (Avg.).
  Res. 0-30% 30-50% 50-70% 70%+ Avg.
BEDLAM 644 65.1 61.0 58.4 64.6 63.3
896 57.8 54.7 54.7 59.3 56.5
1288 55.2 50.4 51.6 56.0 53.2
448* 60.3 60.3 61.0 68.6 60.5
644* 55.6 56.0 57.6 63.1 56.0
896* 51.1 51.3 53.7 58.6 51.4
Res. 0-10% 10-20% 20-30% 30%+ Avg.
AGORA 644 100.8 77.2 59.2 53.0 72.0
896 91.4 71.3 55.3 50.1 67.0
1288 82.2 64.9 52.4 48.2 61.9
448* 94.6 74.1 58.5 54.7 70.0
644* 84.6 68.5 57.3 52.7 65.5
896* 76.5 63.7 54.1 49.9 61.0
Table B3: Dataset comparisons. We report the number of instances for different scale ranges on different datasets, as well as MVE comparisons between our method (with “*”) and the baseline.
  0-30% 30-50% 50-70% 70%+ Avg.
BEDLAM Count 54041 36362 4586 1497 -
644* 55.6 56.0 57.6 63.1 56.0
644 65.1 61.0 58.4 64.6 63.3
3DPW Count 5121 10113 12299 7982 -
644* 86.2 74.2 70.5 70.2 73.7
644 93.6 77.8 68.7 68.3 74.8
0-10% 10-20% 20-30% 30%+ Avg.
AGORA Count 1556 2976 1986 1274 -
644* 84.6 68.5 57.3 52.7 65.5
644 100.8 77.2 59.2 53.0 72.0
0-40% 40%+ Avg.
Panoptic Count 31592 11172 -
644* 85.2 81.0 84.2
644 90.0 82.6 88.2
Table B4: Comparison of different scale thresholds on BEDLAM [3] validation set. We report MVE for different scale ranges, average (Avg.) MVE and inference runtime (ms).
 $\alpha_{\text{s}}$ MVE $\downarrow$ Time (ms)
0-30% 30-50% 50-70% 70%+ Avg.
(a) 0.0 67.0 60.4 58.6 64.0 64.0 37.2
(b) 0.3 60.0 61.9 59.9 65.8 60.8 40.9
(c) 0.5
(Ours) 55.6 56.0 57.6 63.1 56.0 42.0
(d) 0.7 55.9 56.1 58.2 65.2 56.2 44.1
(e) 1.0 58.5 56.8 58.5 63.9 57.9 46.4
@@
Effect of resolution. To study the impact of resolution, we train our single-resolution baseline and our proposed method with scale-adaptive tokens using different resolution settings, reporting estimation errors (MVE) on AGORA [33] validation set and BEDLAM [3] validation set in Tab. B2. As input resolution increases, both single-resolution (baseline) and mixed-resolution (our method, with “*”) settings show accuracy improvements across individuals in different scale ranges. Compared to the corresponding baseline, our method greatly reduces the estimation error in small-scale instances. These results further demonstrate the importance of higher resolution, as it leads to better outcomes, and highlight the effectiveness of our mixed-resolution strategy.
Scale distribution. We report the number of instances for each scale range on AGORA [33] validation set, BEDLAM [3] validation set, 3DPW [49] test set and CMU Panoptic [15] test set in Tab. B3 (MuPoTS [29] is not included due to the lack of bounding box annotations). To study the impact of scale-adaptive tokens, we evaluate our method (644*, mixed resolution) and the corresponding single-resolution baseline (644) on these datasets. As shown, AGORA [33] and BEDLAM [3] contain more small-scale instances, where our method outperforms baseline with reduced MVE. This improvement is also seen on small-scale instances in 3DPW [49] and CMU Panoptic [15]. However, 3DPW’s larger-scale instances show no improvement, likely due to the model’s limited capability.
Scale threshold $\alpha_{\text{s}}$. To further study the impact of high-resolution tokens, we conduct experiments on various scale thresholds $\alpha_{\text{s}}$ while retaining the pooling of background tokens. $\alpha_{\text{s}}=0$ denotes that no high-resolution tokens are used. Results are shown in Tab. B4. Compared to (a), our method (c) achieves a consistent error reduction across different scale ranges, indicating that introducing sufficient high-resolution tokens eases the estimation challenge on small-scale instances and also allows the model to deal better with large-scale instances. In (b), although the improvement in the scale range of 0-30% is significant, the decrease in high-resolution samples increases learning difficulty during training and potentially leads to worse performance on larger-scale instances than (a). (d) and (e) indicate that too large $\alpha_{\text{s}}$ decreases efficiency with longer interence time cost without improving accuracy. In (d), a large scale threshold ignores plenty of background context since the high-resolution tokens are encoded independently for $N_{\text{hr}}$ layers, leading to a performance decline, which is consistent with our findings in Sec. 4.4. In (e), when all humans are processed with high-resolution tokens (i.e. $\alpha_{\text{s}}=1.0$), training becomes unstable and results are worse. In general, $\alpha_{\text{s}}=0.5$ achieves the best speed-accuracy trade-off.
Accuracy and effect of scale map prediction. We evaluate our scale map prediction in Tab. B5 which achieves high prediction accuracy, with 0.98 F1-Score. To further analyze its impact on the final mesh estimation, we replace the predicted scale map with GT scale map during inference time. The average MVE slightly improves, indicating our scale predictions are accurate with minimal impact on accuracy. For scales over 50%, we find that the predicted scale map outperforms GT due to underestimating scales in some cases, assigning more instances of high-resolution tokens, thus improving results.
Table B5: Accuracy and effect of scale prediction on BEDLAM [3] validation set. We report F1-Score (F1) and MAE for evaluating scale map accuracy. To analyze its impact on mesh estimation, we replace the predicted scale map with GT and report MVE for different scale ranges and the average (Avg.).
 Scale F1 $\uparrow$ MAE $\downarrow$ MVE $\downarrow$
map 0-30% 30-50% 50-70% 70%+ Avg.
Pred. 0.98 0.056 55.6 56.0 57.6 63.1 56.0
GT - - 55.4 55.8 58.6 63.2 55.8